OpenStudy (anonymous):

During a sale, a dress decreased in price from $60 to $57. What was the percent decrease ? Simplify your answer

OpenStudy (anonymous):

do u know the formula for percent decrease?

OpenStudy (anonymous):

? do u

OpenStudy (anonymous):

to get %age decrease just divide the new price to the orignal and multiply hundred...then subtract the 100-answer... (57/60)*100=95% now % decrese=(100-95)%=5%

OpenStudy (anonymous):

yeah a quicker formula is

OpenStudy (anonymous):